When Kim Seng bought 8 identical CD he was short $28. When he bought 4 identical CD and 3 tapes he had $17 left. If each tape cost $5 how much money did he have at first.
Thanks

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

Let a = amt of money he had at first
let c = cost of each CD
:
When Kim Seng bought 8 identical CD he was short $28.
a = 8c - 28
When he bought 4 identical CD and 3 tapes he had $17 left.
If each tape cost $5
a = 4c + 3(5) + 17
a = 4c + 15 + 17
a = 4c + 32
a=a, therefore, find c
8c - 28 = 4c + 32
8c - 4c = 32 + 28
4c = 60
c = 60/4
c = $15 for each CD
:
how much money did he have at first?
a = 8(15) - 28
a = 120 - 28
a = $92 amt he had at first
:
:
Check in the 2nd equation
a = 4(15) + 32
a = $92
:

Answer by ikleyn(52778)

.

Just the very first phrase  " When Kim Seng bought 8 identical CD he was short $28 "  makes no sense and is lame.

Indeed,  having short,  Kim Seng  (as anyone else)  would not be able to buy 8 CDs.


Watch and track your writing  ( ! )